# Large Language Model Integration with Reinforcement Learning to Augment Decision-Making in Autonomous Cyber Operations

### v1

@@ Abstract
Reinforcement Learning (RL) has shown great potential for autonomous decision-making in the cybersecurity domain, enabling agents to learn through direct environment interaction. However, RL agents in Autonomous Cyber Operations (ACO) typically learn from scratch, requiring them to execute undesirable actions to learn their consequences. In this study, we integrate external knowledge in the form of a Large Language Model (LLM) pretrained on cybersecurity data that our RL agent can directly leverage to make informed decisions. By guiding initial training with an LLM, we improve baseline performance and reduce the need for exploratory actions with obviously negative outcomes. We evaluate our LLM-integrated approach in a simulated cybersecurity environment, and demonstrate that our guided agent achieves over 2x higher rewards during early training and converges to a favorable policy approximately 4,500 episodes faster than the baseline.
Index Terms:
Autonomous Cyber Operations, Large Language Model, Reinforcement Learning, Cybersecurity, LLM-RL Integration.
@@ I Introduction
The extent to which offensive cyber operations have increased over recent years is significant [1, 2]. To counter this trend, cybersecurity must scale proportionally and sustainably. Autonomous Cyber Operations (ACO) enable agents to make decisions on behalf of human operators, significantly enhancing the scalability of cybersecurity [3]. Modern ACO applications are made possible through Reinforcement Learning (RL), where agents learn through direct environment interaction [4]. However, current RL applications in ACO require these agents to learn from scratch, leading to agents that must:
  • Perform undesirable actions to learn their associated consequences; and
  • Take longer to converge onto a favorable policy [5, 6, 7].
We address these limitations by incorporating external knowledge in the form of a Large Language Model (LLM) into the RL pipeline, which the agent can directly leverage to augment its decision-making capabilities. We validate our approach using CybORG, a widely used cybersecurity environment designed to train blue agents to defend an enterprise network against adversarial threats [3].
@@
  • The Methodology section describes the design of our research activities and experimental setup.
  • The Evaluation section presents our results and explains the rationale behind key implementation decisions.
@@
This paper extracts work from the author’s thesis, submitted to the Royal Military College of Canada (RMC) [2].
@@ II-A LLMs and RL
where $\mathcal{L}_{RL}(\theta)$ is the standard PPO loss, and
$\mathbb{E}_{s\sim\pi_{\theta}}\mathcal{H}\left(\pi_{T}(\cdot|s)\|\pi_{\theta}(\cdot|s)\right)$ represents the deviation between the RL agent’s and the teacher’s policy. The weighting coefficient $\lambda$ is gradually decayed, reducing the LLM’s influence over time.
@@ III-A LLM Evaluation
TABLE I: LLMs involved during the evaluation process
LLM Name Baseline Architecture LLM Source
@@ III-C3 Integrating the LLM
TABLE II: Final hyperparameters used for the LLM integration
@@ IV-A LLM Evaluation
As discussed in Section III-A, we evaluated the four LLMs outlined in Table I against a dataset of 100 questions using the prompt format shown in Fig. 1. To improve the contextual relevance of the LLMs’ responses, we formatted CybORG’s raw output into a condensed JSON and sentence format. We compare the results using the raw versus transformed inputs in Appendix B.
@@
TABLE III: Limitations of relying solely on BERTScore for evaluating LLM responses
Best Action (Validation Label) LLM Response Precision
Restore Operation1 Allow Enterprise1 80 0.8675
@@ IV-B1 Method Summary
TABLE IV: Manual scoring of each LLM across 20 easy, 40 medium, and 40 hard questions in both JSON and sentence formats. The highest scores per row are bolded
@@
Easy Scores (JSON) Total 8 4.5 12 9.5
@@
Medium Scores (JSON) Total 7 10 19.5 14
@@
Hard Scores (JSON) Total 4 7.5 14 12
@@
Easy Scores (Sentence) Total 1 6 6 9
@@
Medium Scores (Sentence) Total 2.5 16 8 13.5
@@
Hard Scores (Sentence) Total 1 13 4 13
@@
Total Scores (JSON) Total 19 22 45.5 35.5
@@
Total Scores (Sentence) Total 4.5 35 18 35.5
@@ IV-B2 Performance Analysis
TABLE VI: LIME results for feature space modification with action masking. Only the weight of the one-hot encoded feature is shown. ’Reco in Top 4’ indicates whether the teacher’s recommendation is within the top four actions of the RL agent’s policy, with its associated ranking
@@ IV-C1 Prompt Design
Fig. 7 illustrates the performance of the four prompt formats we tested for the LLM integration. Prompt 4, denoted by the green curve, was ultimately used for our study due to its superior performance and stability.
@@ IV-C3 LLM Guidance as a Distribution
TABLE VII: Comparison of the top three action probabilities for the LLM-guided agent, and agents trained independently for 100 and 10,000 episodes. Column titles are abbreviated for readability
@@
We describe the procedure for mapping the LLM’s output into a probability distribution and using this to guide training in Appendix C.
@@ Appendix C Mapping the LLM to a Distribution
We present the results of incorporating the LLM’s guidance as a distribution, instead of a single action, in Fig. 9 under Section IV-C.

### v2

@@ Abstract
Reinforcement Learning (RL) has emerged as a promising approach for autonomous decision-making in the cybersecurity domain, enabling agents to learn through direct environment interaction. However, RL agents in Autonomous Cyber Operations (ACO) typically learn from scratch, requiring them to execute undesirable actions to learn their consequences. In this study, we integrate external knowledge in the form of a Large Language Model (LLM) pretrained on cybersecurity data that our RL agent can directly leverage during training to make informed decisions. By guiding initial training with an LLM, we improve baseline performance and reduce the need for exploratory actions with clearly negative outcomes. We evaluate our LLM-integrated approach in a realistic, simulated cybersecurity environment, and demonstrate that our guided agent achieves over 2x higher rewards during early training and converges to a favorable policy approximately 4,500 episodes faster than the baseline.
@@ I Introduction
Offensive cyber operations have increased at an unprecedented scale over recent years [17, 35, 6, 8]. It is no longer feasible for humans to manually defend their IT systems, motivating the development of automated tools to help with the intrusion detection process [10]. These tools began as signature-based approaches, which require known attacks to exist in a database to be successfully detected [31]. Because an exact match is required for detection, these systems are vulnerable to any new exploits whose signatures are not yet stored.
Machine Learning (ML) approaches helped overcome these deficiencies by training models to detect patterns across malicious and benign samples, eliminating the requirement for exact matches [9]. However, traditional supervised approaches depend on large, static datasets to train models. These datasets must be manually updated whenever new behaviors emerge; a task that is not scalable given the rapidly evolving nature of cybersecurity.
Autonomous Cyber Operations (ACO) is an emerging field in which specialized cybersecurity agents are trained to make effective decisions on behalf of humans [3]. Unlike traditional ML, ACO employs Reinforcement Learning (RL), where agents can be trained through direct interaction with their environment, eliminating the reliance on large, static, labeled datasets [34]. This interaction enables agents to dynamically adapt to evolving adversarial Tactics Techniques and Procedures (TTPs), offering a scalable and adaptive approach to modern cyber defense.
For ACO to operate effectively, an environment must exist that mimics cybersecurity, and produces the necessary signals for an agent to learn. The Technical Cooperation Program (TTCP) has created such an environment: CybORG, a widely used RL framework designed to train agents to defend a simulated enterprise network against adversarial threats [3, 19].
Despite this progress, RL agents in ACO typically begin as “bare-bone” models, requiring substantial training time before converging to an optimal policy [43, 24]. Moreover, because the agents’ weights are initially randomized, they are equally likely to select favorable or unfavorable actions. This poses significant risks in the context of cybersecurity, where an incorrect action can lead to the compromise of entire IT systems.
In addition, the data involved in the incident detection and response process is typically textual by nature (e.g., log files) and must be mapped to a numerical feature space for RL agents to process. This feature engineering process is typically deterministic, and could miss critical information relating to cybersecurity incidents [3].
These limitations can be mitigated by incorporating external knowledge into the RL pipeline that agents can directly leverage during decision-making. Instead of acting randomly and learning solely from environmental feedback, an agent can benefit from a teacher that provides guidance, allowing continued learning from a strong baseline [28]. Such a teacher should be generalizable across cybersecurity environments, and capable of effectively identifying patterns in the textual information present in cybersecurity data.
In this study, we integrate a Large Language Model (LLM) into the ACO pipeline to improve an RL agent’s decision-making capabilities. The vast amounts of textual data used to train LLMs make them ideal candidates to serve as a teacher, enabling them to generalize across cybersecurity environments and capture contextually relevant information that might otherwise be missed during feature engineering [46].
@@
  • The Methodology section describes the design and implementation of our research activities and experimental setup.
  • The Evaluation section presents our results and explains the rationale behind major implementation decisions.
@@ II-A LLMs and RL
where $\mathcal{L}_{RL}(\theta)$ is the standard PPO loss, and the term $\mathbb{E}_{s\sim\pi_{\theta}}\mathcal{H}\left(\pi_{T}(\cdot|s)\|\pi_{\theta}(\cdot|s)\right)$ represents the deviation between the RL agent’s and the teacher’s policy. The weighting coefficient $\lambda$ is gradually decayed, reducing the LLM’s influence over time.
@@ III-A LLM Evaluation
TABLE I: LLMs involved during the evaluation process.
LLM Name Baseline Architecture Source
@@ III-C3 Integrating the LLM
III-D Evaluation Design
III-D1 Quantifying the Teacher’s Influence
TABLE II: Final hyperparameters used for the LLM integration.
@@ IV-A LLM Evaluation
As discussed in Section III-A, we evaluated the four LLMs outlined in Table I against a dataset of 100 questions using the prompt format shown in Fig. 1. To improve the contextual relevance of the LLMs’ responses, we formatted CybORG’s raw output into a condensed JSON and sentence format. We compare the results using the raw versus transformed inputs in Appendix B of the online supplemental material.
@@ IV-B1 Method Summary
TABLE III: Limitations of relying solely on BERTScore for evaluating LLM responses.
Best Action (Validation Label) LLM Response Precision
Restore Operation1 Allow Enterprise1 80 0.8675
TABLE IV: Manual scoring of each LLM across 20 easy, 40 medium, and 40 hard questions in both JSON and sentence formats. The highest scores per row are bolded.
@@
Easy (JSON) Total 8 4.5 12 9.5
@@
Medium (JSON) Total 7 10 19.5 14
@@
Hard (JSON) Total 4 7.5 14 12
@@
Easy (Sentence) Total 1 6 6 9
@@
Medium (Sentence) Total 2.5 16 8 13.5
@@
Hard (Sentence) Total 1 13 4 13
@@
Total (JSON) Total 19 22 45.5 35.5
@@
Total (Sentence) Total 4.5 35 18 35.5
@@ IV-B2 Performance Analysis
TABLE VI: LIME results for feature space modification with action masking. Only the weight of the one-hot encoded feature is shown. ’Reco in Top 4’ indicates whether the teacher’s recommendation is within the top four actions of the RL agent’s policy, with its associated ranking.
@@ IV-C1 Prompt Design
Fig. 7 illustrates the performance of the four prompt formats we tested for the LLM integration. Prompt 4, denoted by the green curve, was ultimately used for our study.
@@ IV-C3 LLM Guidance as a Distribution
TABLE VII: Comparison of the top three action probabilities for the LLM-guided agent, and agents trained independently for 100 and 10,000 episodes. Column titles are abbreviated for readability.
@@
We describe the procedure for mapping the LLM’s output into a probability distribution and using this to guide training in Appendix C of the online supplemental material.
@@ Appendix C Mapping the LLM to a Distribution
We present the results of incorporating the LLM’s guidance as a distribution in Fig. 9 under Section IV-C.